What is the mass in grams of CuO that would be required to produce 41.50 moles of nitrogen gas in the following chemical reaction?

2 NH₃(g) + 3 CuO(s) → 3 Cu(s) + N₂(g) + 3 H₂O(g)



1
Expert's answer
2021-02-22T05:49:34-0500

From the balanced chemical,

Moles of CuO required = 3×41.5 = 124.5 mol

Mass of CuO = 124.5 × 56 = 6972 g

Mass of CuO =[ 6972 ] g